# When Does Multi-Task Learning Fail? Quantifying Data Imbalance and Task Independence in Metal Alloy Property Prediction

### v1

@@ 5.4.4 Soft Parameter Sharing
Our MTL architectures used hard parameter sharing (fully shared backbone). Soft parameter sharing approaches, which allow different tasks to have different parameters with regularization encouraging similarity, may be more robust to task dissimilarity. This remains an avenue for future investigation.
@@ 6 Conclusion
This study critically examined the effectiveness of multi-task learning for metal alloy property prediction, challenging the common assumption that MTL universally improves performance when properties share underlying physics.
@@
Future work should investigate gradient balancing methods, curriculum learning approaches, and soft parameter sharing architectures to potentially unlock the benefits of MTL even when tasks are weakly related. Additionally, expanding this analysis to other material systems and property combinations will help establish when MTL is truly beneficial in materials discovery.

### v2

@@ 5.5.4 Soft Parameter Sharing
Our soft parameter sharing experiments (Section 4, Experiment 6) reveal that allowing task-specific backbone deviations provides complementary benefits to gradient-based methods:
  • 1.
    Minority task improvements: With $\lambda=0.001$, hardness $R^{2}$ improved from 0.735 to 0.764 (+3.9%) and amorphous F1 improved from 0.714 to 0.744 (+4.2%), at a small cost to resistivity (-0.8%).
  • 2.
    Tail performance: Soft sharing reduced the maximum hardness prediction error from 3.11 to 1.95 (37% reduction), indicating improved robustness for extreme values.
  • 3.
    Backbone similarity: Despite separate backbones, the learned similarity remained high (99.6%), indicating that only subtle task-specific adaptations are needed.
Soft parameter sharing provides a trade-off compared to gradient-based methods: while PCGrad achieves higher overall hardness $R^{2}$ (0.855 vs. 0.764), soft sharing may be preferred when tail performance is critical or when gradient manipulation introduces undesirable instabilities.
5.5.5 Feature Distribution Smoothing
FDS (Section 4, Experiment 5) showed no measurable improvement in our setting. This negative result is informative: unlike natural image datasets where FDS was originally developed, alloy composition features are already well-distributed in the compositional simplex. The feature-space imbalance addressed by FDS is distinct from the label-space and task-level imbalance that drives negative transfer in our materials informatics setting. This suggests that DIR techniques should be selected based on the specific nature of the imbalance—feature-space smoothing is ineffective when the imbalance is primarily at the task or label level.
@@ 6 Conclusion
This study critically examined the effectiveness of multi-task learning for metal alloy property prediction, challenging the common assumption that MTL universally improves performance when properties share underlying physics. By integrating our empirical findings with the Deep Imbalanced Regression (DIR) framework [yang2021delving], we provide both practical guidelines and theoretical insight into the mechanisms of negative transfer.
@@
  5. 5.
    DIR mitigation techniques effectively address negative transfer. PCGrad improved hardness $R^{2}$ by 12.4% (0.761 $\rightarrow$ 0.855), confirming that gradient conflicts were the primary bottleneck. LDS+GradNorm achieved the best overall balance with improvements across all tasks. However, combining techniques does not guarantee additive benefits—LDS+PCGrad showed negative synergy.
@@
  • 5.
    Apply MTL within property clusters (electronic, mechanical, thermodynamic) rather than across them
  • 6.
    Use PCGrad when maximizing minority-task performance is the priority
  • 7.
    Use LDS+GradNorm when balanced multi-task performance is desired
  • 8.
    Do not assume combining mitigation techniques yields additive benefits
We propose a “materials taskonomy” framework that organizes properties by their underlying physics. Properties within the same cluster—such as hardness and yield strength (mechanical) or resistivity and thermal conductivity (electronic)—are expected to benefit from joint learning, while cross-cluster MTL requires careful validation.
Our evaluation of DIR techniques—including LDS, Balanced MSE, PCGrad, and GradNorm—demonstrates that gradient conflict resolution can substantially mitigate negative transfer, with PCGrad achieving +12.4% improvement on the minority task. Future work should explore active learning approaches [rao2022machine] that curate informative training distributions, and sequential transfer learning [kamnis2024transformer] where models are pre-trained on large datasets and fine-tuned for specific properties.
@@
Data Availability
The metal alloy property datasets used in this study are available from the AI-Hub Korea platform [aihub2022material]. Access requires free registration. The datasets include: (1) electrical resistivity measurements (52,388 samples), (2) Vickers hardness measurements (800 samples), and (3) amorphous-forming ability labels (840 samples). All experimental code, including implementations of LDS, FDS, PCGrad, GradNorm, and soft parameter sharing, is available at https://github.com/[repository] under the MIT license.
CRediT Author Statement
Author One: Conceptualization, Methodology, Software, Writing - Original Draft. Author Two: Validation, Investigation, Writing - Review & Editing. Author Three: Supervision, Project Administration, Writing & Editing.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This research was supported by [Funding Agency]. The authors thank [acknowledgments].